On July 31, 2017, Amsterdam Company engaged Minsk Tooling Company to construct a special- purpose piece of factory machinery. Construction was begun immediately and was completed on November 30, 2017. To help finance construction, on July 31 Amsterdam issued a $80,000, 3 year, 11% notes payable, on which interest is payable each July 31. In addition, the company had outstanding all year a 8%, 3-year, $100,000 note payable and an 10%, 4-year, $150,000 note payable. Interests on both 8% and 10% note payables are payable each December 31. Amsterdam made payments to Minsk of

$240,000 on July 31 and $120,000 on October 1.

  1. Calculate weighted-average accumulated expenditures avoidable interest, and total interest cost to be capitalized during 2017.
  2. Prepare the journal entries needed on the books of Amsterdam Company at each of the following dates (July 31, October 1, and December 31).
Answer #1

a.

weighted-average accumulated expenditures avoidable interest, and total interest cost to be capitalized during 2017.

Date price time wighted Avg. Cost
July 31,2017 $240000 4/12 $240000x4/12 =$80000
October 1,2017 $120000 2/12 $120000x2/12 = $20000
Total $100000
Total Contract cost $360000
Total Borrowing $330000
Funded By Own Capital $30000

Specific Interest Rate = 11%

Generic Borrowing-

price time Outstanding Weighted average borrowing Interest
8%notes payable $100000 12 month $100000 $8000
10% Notes payable $150000 12 Month $150000 $15000
Total $250000 $23000

Weighted avg. Generic Borrowing inst rate = $23000/$250000 * 100 = 9.2%

-removed interest is the interest on the particular Borrowing only. where the interest expenses on the generic taking will be incurred irrespective of hence the Construction will be set or Not.

-Avoidable interest = $80000*11%*5/12  = $3667

-estimation on Interest Capitalization

Date price Amount Allocated To general Borrowing Weighted for the period Outstanding
July 31 $240000

$160000

(balance $80000 will be given from specific Borrowing) on $80000 notes payment)

$160000x4/12 = $53333
October 1 $120000

$90000

(Balance $30000 will be given from own capital)

$90000x2/12 = $15000
Total $360000 $250000 $68333

Borrowing cost to be capitalised-

On Specific Borrowing    =$80000x11%x4/12 = $2933   
On Generic Borrowing =$68333x9.2% = $6287
Total $9220

-------------------------

(b)

Journal entry -

Date Accounts Debit Credit
July 31 Cash $80000
11% Notes payment $80000
july 31 Machinery (Wok in process) 240000
Cash 240000
Oct 1 Machinery (Work In process) 120000
Cash 120000
Nov 30 Machinery 360000
Machinery (Work in progress) 360000
Dec31 Machinery $9220
Interest Expenses $17447
Interest payment(8% Notes) $8000
Interest payment(10% Notes) $15000
Interest payment(11% Notes) $3667
Dec31 interest payment $8000
Interest payment(10% Notes) $15000
Cash $23000